OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = (x - 3)/2, what is the value of f[g(-5)]? = f[-4] = 2(-4) + 3 = -8 + 3 = -5

OpenStudy (anonymous):

seems right

OpenStudy (anonymous):

are you sure lol

OpenStudy (anonymous):

hang on while i do the problem

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = (x - 3)/2, what is the value of f[g(-5)]? g(-5)=(-5-3)/2=-4 f(g(-5))=f(-4)=2*(-4)+3=-8+3=-5 So yeh -5

OpenStudy (anonymous):

okay thanks!